“Gasohol” is a mixture of gasoline and ethyl alcohol used as automobile fuel. The alcohol produces energy in a combustion reaction with O2.
C2H5OH(l) + 3O2(g) → 2CO2(g) + 3H2O(l)
If 0.115 g of alcohol evolves 3.42 kJ when burned at constant pressure, what is the molar enthalpy of combustion for ethyl alcohol?
1
Expert's answer
2015-11-02T12:26:24-0500
The enthalpy of combustion of ethyl alcohol will be the heat divided by number of the moles. The number of the moles can be calculated from the mass of the ethanol and molar mass as n= m/M. The molar mass of ethyl alcohol is 46.07 g/mol. Thus, one has an enthalpy of combustion:
∆Hc = Q/n = Q/m/M = 3.42kJ/0.115g/46.07g/mol = 1370kJ/mol